OpenStudy (anonymous):

Match each percent problem with its solution. 1. 10% of what number is 70? 2. 7 is what percent of ten? 3. What is 70% of 100? 4. 7 is 7% of what number? a. 700 b. 70 c. 98 d. 100

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

percent formula : is/of = %/100 1. 10% of what number is 70 % = 10 of = x is = 70 now sub into the formula 70/x = 10/100 cross multiply 10x = 7000 x = 700 2. 7 is what percent of 10 70% 3. what is 70% of 100 .7(100) = 70 4. 7 is 7% of what number 7 = .07x 7/.07 = x 100 = x